need it asap!!!!!

thank you!!

on January 1 Ogden inc had $1500 of supplies on hand. during January, Odgen purchased $5500 worth of new supplies. at the end of the month, a count revealed $700 worth of supplies remaining on the shelves. prepare the adjusting journal entry required at the end of the month. included date ...omit explanations.
Answer #1

Given:

Opening Balance = $1500

Purchased Supplies= $5500

Closing Balance = $700

Closing Balance = Opening Balance + Purchased Supplies - Supplies used

700=1500+5500-Supplies used

Supplies used = 7000-700=$6300

Adjusting Entry

Date Particulars Amount($) Amount($)
Jan 31 Supplies Expenses 6300
Supplies 6300

Note:

The used supplies are recorded as an expenses for the period. Thus, the supplies (as an assets account) decrease whereas the used supplies (supplies expenses account) increase.